### v2

# Anytime-valid simultaneous lower confidence bounds for the true discovery proportion

Abstract
We propose a method that combines the closed testing framework with the concept of safe anytime-valid inference (SAVI) to compute lower confidence bounds for the true discovery proportion in a multiple testing setting. The proposed procedure provides confidence bounds that are valid at every observation time point and that are simultaneous for all possible subsets of hypotheses. While the hypotheses are assumed to be fixed over time, the subsets of interest may vary. Anytime-valid simultaneous confidence bounds allow us to sequentially update the bounds over time and allow for optional stopping. This is a desirable property in practical applications such as neuroscience, where data acquisition is costly and time-consuming. We also present a computational shortcut which makes the application of the proposed procedure feasible when the number of hypotheses under consideration is large. We illustrate the performance of the proposed method in a simulation study and give some practical guidelines on the implementation of the proposed procedure.
$\mathbf{Keywords}$: multiple testing, selective inference, sequential testing, e-processes
1 Introduction
When testing multiple hypotheses, researchers might be interested in grouping hypotheses and making inference on these groups of hypotheses. Confidence bounds for the number of false null hypotheses (called true discoveries) in a (sub-)set of considered null hypotheses can be utilized in this case.
If the subset of null hypotheses of interest is specified a-priori, procedures to compute lower confidence bounds for the true discovery proportion (TDP) are given in Meinshausen and Rice (2006); Ge and Li (2012); Patra and Sen (2016); Preusse et al. (2025), among others. Simultaneous lower confidence bounds for the TDP make inference on the true discoveries simultaneously for all possible hypothesis (sub-)sets. This allows the practitioner to specify sets of hypotheses of interest based on the corresponding confidence bounds for the TDP. Methods to compute simultaneous lower confidence bounds for the TDP have been proposed by Genovese and Wasserman (2006); Goeman and Solari (2011); Hemerik et al. (2019); Blanchard et al. (2020); Katsevich and Ramdas (2020); Vesely et al. (2023), among others, and have been applied to genomics (Ebrahimpoor et al., 2020) or functional magnetic resonance imaging (fMRI) (Rosenblatt et al., 2018; Blain et al., 2022; Andreella et al., 2023). The aforementioned procedures are based on closed testing (Marcus et al., 1976), which has been shown to be the only admissible framework for confidence bounds for the TDP (Goeman et al., 2021b).
These established methods require a fixed sample size; early stopping or increasing the sample size after the data analysis is not allowed. Time and/ or monetary restraints might require stopping the data acquisition- and analyze the available data- earlier than planned. Extensions of deadlines or additional funding might lift these constraints at a later time point and being able to resume the study seems advantageous. Even if there are no restraints, stopping the data gathering process as soon as the results are deemed to be interesting can save both time and money compared to sampling until a pre-specified sample size is achieved. Deciding based on the current results or based on external rules whether data acquisition should be continued is called optional stopping (c.f. Grünwald et al., 2024). Even if the data acquisition is stopped and the data are analyzed, optional stopping allows for the continuation of the data acquisition at a later time point and combining the new data with the previous results. Procedures that are valid under optional stopping are often called anytime-valid procedures.
Sequentially testing a single hypothesis under optional stopping can be done using an e-process, which is an extension of the recently proposed e-value (see e.g. Vovk and Wang, 2021; Grünwald et al., 2024; Koolen and Grünwald, 2022; Ruf et al., 2023). An e-value can be interpreted as measuring the evidence against a null hypothesis (Vovk and Wang, 2021). An e-process is a process of anytime-valid e-values, that is, e-values that are valid at any time point of the process. The construction of e-processes is an active topic of research (Koolen and Grünwald, 2022; Ramdas et al., 2023; Pérez-Ortiz et al., 2024; ter Schure et al., 2024; Turner et al., 2024; Vovk and Wang, 2024a; Wang and Ramdas, 2025, among others).
Simultaneous lower confidence bounds for the TDP based on e-values have been proposed by Vovk and Wang (2023, 2024b). Their approach does not allow for updating the confidence bounds over time. Fischer and Ramdas (2025) update the lower confidence bounds for the TDP using e-values. In contrast to our work, they focus on the online setting, that is, the hypotheses of interest are updated over time. We study the offline setting, where the hypotheses of interest are considered to be fixed.
In this work, we propose a procedure to compute anytime-valid simultaneous lower confidence bounds for the TDP in the offline setting. To this end, we combine the closed testing framework with e-processes. Furthermore, we derive a computational shortcut so that computation of the bounds is feasible if the number of hypotheses of interest is large. The novelty of our procedure is that it allows for optional stopping of the data acquisition.
The rest of the work is structured as follows. In the next section, we specify our framework and review related literature, focusing on simultaneous confidence bounds for the TDP based on closed testing as well as e-values and e-processes. We propose the procedure used to compute anytime-valid simultaneous confidence bounds for the TDP in Section 3. In Section 4, we discuss some practical considerations. We investigate the behavior of the proposed method in a simulation study in Section 5 and end with a short discussion in Section 6.
2 Notation and Related Literature
First, we introduce the notation and then we give a brief review of simultaneous confidence bounds for the TDP based on closed testing. We end this section with a short overview of e-values and e-processes.
In the following, we use bold letters to refer to vectors or, if explicitly stated, matrices. Furthermore, when we write hypothesis we are referring to a null hypothesis, unless explicitly stated otherwise.
We consider the probability space $(\Omega,\mathcal{F},\mathcal{P})$, where $\Omega$ is a set of vectors of length $m$ and $\mathcal{P}$ is the set of all probability measures on $(\Omega,\mathcal{F})$. A probability distribution is denoted by $\mathbb{P}$ with corresponding expected value $\mathbb{E}_{\mathbb{P}}$. The probability measure under the true but unknown data generating process is denoted by $\mathbb{P}^{\ast}$, $\mathbb{P}^{\ast}\in\mathcal{P}$.
We aim to make inference on the set of hypotheses $\mathcal{H}\coloneq\{H_{1},\ldots,H_{m}\}$. We refer to $H_{i}$, $i\in\{1\ldots,m\}$, as elementary hypothesis. The considered hypotheses are composite hypotheses, that is, each elementary hypothesis $H_{i}$ corresponds to a set of probability measures $\mathcal{P}_{0}(i)\subset\mathcal{P}$. The hypothesis $H_{i}$ is true if and only if $\mathbb{P}^{\ast}\in\mathcal{P}_{0}(i)$.
Let $\mathcal{C}=2^{\{1,\ldots,m\}}\setminus\emptyset$ denote the collection of all nonempty subsets of $\{1,\ldots,m\}$. Intersection hypotheses are denoted by $H_{I}\coloneq\bigcap_{i\in I}H_{i}$, where $I\in\mathcal{C}$ is the set of indices corresponding to the elementary hypotheses in $H_{I}$. The size of set $I$ is denoted by $|I|$, with $|I|\geq 1$. Therefore, $H_{I}$ might refer to an elementary or intersection hypothesis, while $H_{i}$ always denotes an elementary hypothesis. We denote by $\mathcal{P}_{0}(I)\coloneq\bigcap_{i\in I}\mathcal{P}_{0}(i)$ the set of all probability measures on $(\Omega,\mathcal{F})$ corresponding to $H_{I}$. The set $T\subseteq\{1,\ldots,m\}$ refers to the index set of all true elementary hypotheses, which implies that $\mathbb{P}^{\ast}\in\mathcal{P}_{0}(T)$. Since $H_{T}$ includes all true elementary hypotheses, $\mathbb{P}^{\ast}\notin\mathcal{P}_{0}(i)$ for every $i\notin T$. Thus, an intersection hypothesis $H_{I}$ is true if and only if $I\subseteq T$.
The set of indices of the rejected elementary hypotheses is denoted by $R$, with $R\in\mathcal{C}$. Note that we do not specify any rejection rule to determine $R$. Indeed, the rejected hypotheses correspond to (a subset of) the hypotheses that are of interest to the practitioner. Every rejected hypothesis is considered to be a discovery and a true hypothesis corresponds to a false discovery. Because $R$ specifies the set of discoveries, we call it a discovery set. Denote by $T_{R}=T\cap R$ the set of indices corresponding to true elementary hypotheses in the discovery set. The number of false discoveries is denoted by $|T_{R}|=\tau(R)$. The false discovery proportion (FDP) is given by $\pi_{0}(R)=\tau(R)/|R|$ and the true discovery proportion (TDP) is given by $\pi_{1}(R)=1-\pi_{0}(R)$.
Lastly, let $\phi_{\alpha}$ denote a family of local level-$\alpha$ test which rejects hypothesis $H_{I}$ iff $\phi_{\alpha}(H_{I})=1$ and which controls the type I error rate at level $\alpha$, i.e., $\mathbb{P}(\phi_{\alpha}(H_{I})=1)\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}_{0}(I)$ and any $\alpha\in(0,1)$.
2.1 Simultaneous lower confidence bounds for the TDP
In this section, we provide more details about the construction of simultaneous lower confidence bounds for $\pi_{1}(R)$. Since $|R|$ is observable, $\tau(R)$ is the only unknown quantity in $\pi_{0}(R)$ and $\pi_{1}(R)$. Thus, to compute lower confidence bounds for $\pi_{1}(R)$, it is sufficient to compute upper confidence bounds for $\tau(R)$. Genovese and Wasserman (2006) and Goeman and Solari (2011) propose simultaneous upper confidence bounds for $\tau(R)$ based on closed testing. Goeman et al. (2021b) show that methods to compute confidence bounds for $\tau(R)$ are either based on closed testing or uniformly improved by closed testing. Closed testing is a class of multiple testing procedures proposed by Marcus et al. (1976). It can be used to determine a rejection set while controlling the family wise error rate, i.e., controlling the probability of making at least one false rejection. Denote by $\mathcal{X}_{\alpha}$ the index set of the hypotheses rejected by closed testing. Let $\mathcal{U}_{\alpha}$ be the index set of the hypotheses that are rejected by a given local level-$\alpha$ test $\phi_{\alpha}$. The intersection hypothesis $H_{I}$ is rejected by closed testing based on $\phi_{\alpha}$, i.e., $I\in\mathcal{X}_{\alpha}$, if and only if for every $J\supseteq I$, $J\in\mathcal{U}_{\alpha}$. Goeman and Solari (2011) showed that
$c_{\alpha}(R)\coloneq\max\{|I|:I\in\mathcal{C}_{R},I\notin\mathcal{X}_{\alpha}\}\geq\tau(R)\quad\text{with }\mathcal{C}_{R}\coloneq\{I\in\mathcal{C}:I\subseteq R\}$ (1)
is a simultaneous $1-\alpha$ upper confidence bound for $\tau(R)$, i.e.,
$\mathbb{P}\left(\tau(R)\leq c_{\alpha}(R)\quad\forall R\in\mathcal{C}\right)\geq 1-\alpha,\qquad\text{for all }\mathbb{P}\in\mathcal{P}_{0}(T)\text{ and }\alpha\in(0,1).$ (2)
Equation (2) holds because $\mathbb{P}(T\notin\mathcal{U}_{\alpha})\geq 1-\alpha$ for all $\mathbb{P}\in\mathcal{P}_{0}(T)$ and therefore $\mathbb{P}(T_{R}\notin\mathcal{X}_{\alpha})\geq 1-\alpha$ for all $\mathbb{P}\in\mathcal{P}_{0}(T)$ and all $R\in\mathcal{C}$. In other words, the probability that $H_{T_{R}}$ is rejected by closed testing is less than $\alpha$, since this event requires that $H_{T}$ is rejected by $\phi_{\alpha}$, which has a probability of less than $\alpha$. Because the event $T\notin\mathcal{U}_{\alpha}$ does not depend on the discovery set $R$, the bounds are simultaneously valid for all $R\in\mathcal{C}$.
This approach requires testing up to $2^{m}-1$ hypotheses. When the local tests are based on p-values, Goeman et al. (2019) propose a shortcut which reduces the number of hypotheses that need to be tested explicitly.
This procedure to compute simultaneous upper confidence bounds for $\tau(R)$ has been extended in recent years, using for example sum tests (Vesely et al., 2023) or permutation tests (Hemerik et al., 2019) as local level-$\alpha$ tests. The approach has also been adapted to different settings, such as knock-offs (Li et al., 2024). Vovk and Wang (2023) extend the approach by Goeman and Solari (2011) to e-values, which are discussed in the following.
2.2 E-variables and e-processes
First, we give a formal definition of an e-variable.
Definition 1
An e-variable $E_{I}$ corresponding to hypothesis $H_{I}$ is a nonnegative random variable with $\mathbb{E}_{\mathbb{P}}[E_{I}]\leq 1$ for all $\mathbb{P}\in\mathcal{P}_{0}(I)$.
The realization of $E_{I}$ is called an e-value, we denote it by $e_{I}$. Because of Markov’s inequality, $\mathbb{P}(E_{I}\geq 1/\alpha)\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}_{0}(I)$ and $\alpha\in(0,1)$. A local level-$\alpha$ test is thus given by $\phi_{\alpha}^{[E]}(H_{I})\coloneq\mathbb{I}\{e_{I}\geq 1/\alpha\}$, where $\mathbb{I}(\cdot)$ denotes the indicator function (Grünwald et al., 2024). It is possible to calibrate e-values based on p-values and vice-versa. Most notable, $P_{I}^{(E)}\coloneq\min(1/E_{I},1)$ is a valid, albeit conservative p-variable (Grünwald et al., 2024). However, the inverse of a p-variable does not need to be a valid e-variable. Data can be analyzed using both e-values and p-values, this has been investigated by Ignatiadis et al. (2024). There are many different families of e-variables. Most often, they are based on a likelihood ratio or Bayes factor (c.f. Grünwald et al., 2024; Wasserman et al., 2020). There are some guidelines on how to choose an e-variable in a specific setting. Mainly, the e-variable should grow as fast as possible if the evidence against the corresponding null hypothesis increases. Grünwald et al. (2024) introduce the growth-rate optimal in the worst-case (GROW) criterion. This criterion measures how fast an e-variable can accumulate evidence against the null hypothesis under the worst-case alternative. It can be used to compare families of e-variables or specify e-variables.
In contrast to p-variables, merging e-variables is rather straight forward. Functions to merge e-variables, so called e-merging functions, have been studied by Vovk and Wang (2021). The e-variable corresponding to the intersection hypothesis $H_{I}$ can be computed using the average as e-merging function, such that $E_{I}=\sum_{i\in I}E_{i}/|I|$. Under arbitrary dependence between the e-variables, the average is the only admissible e-merging function (Vovk and Wang, 2021), meaning that there exists no other e-merging function that returns valid e-variables larger than the e-variable returned by the average. This statement does not hold true for independent e-variables. In this case, the product function is a useful e-merging function (Vovk and Wang, 2021).
As mentioned above, simultaneous lower confidence bounds for the TDP using closed testing can be based on a local level-$\alpha$ e-test. Vovk and Wang (2023) propose discovery e-vectors for the number of true discoveries in a discovery set $R$ based on e-variables with arbitrary dependence. For each $j=1,\ldots,|R|$, the discovery e-vector is given by
$\mathbf{D}^{R}_{F}[j]\coloneq\min_{I\subseteq\{1,\ldots,m\}}F\left((E_{i})_{i\in I}\right),$
where $F$ is a symmetric e-merging function, such as the average. Denote by $\mathbf{d}^{R}_{F}$ the observed discovery e-vector. A discovery e-vector has the following interpretation: the number of true discoveries exceeds $j$ with a probability of at least $1-1/\mathbf{d}^{R}_{F}[j]$. These discovery e-vectors are the inverse of the confidence bounds by Goeman and Solari (2011). If $\mathbf{d}^{R}_{F}[j]\geq 1/\alpha$, then $c_{\alpha}(R)\leq|R|-j$. See Vovk and Wang (2023) for a detailed comparison between discovery e-vectors and the approach by Goeman and Solari (2011). Vovk and Wang (2024b) refine discovery e-vectors for independent and sequential e-values. A sequential e-variable is defined as a random variable with $\mathbb{E}_{\mathbb{P}}[E_{i}|E_{1},\ldots,E_{i-1}]\leq 1$ for all $\mathbb{P}\in\mathcal{P}_{0}(i)$. Sequential e-values can be utilized for temporal processes when the hypotheses are updated with time.
In our approach, we are interested in making inference on a set of hypotheses, which is considered to be fixed, at multiple observation time points, while allowing for optional stopping. This can be done using e-processes. Let $(\mathcal{F}_{n})_{n\geq 0}$ be a filtration, i.e., a sequence of sigma-algebras $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\ldots\subset\mathcal{F}$.
Definition 2
An e-process corresponding to $H_{I}$ is a nonnegative process $(E_{I}^{[n]})_{n\geq 0}$ adapted to some filtration $(\mathcal{F}_{n})_{n\geq 0}$ with $\mathbb{E}_{\mathbb{P}}[E^{[\nu]}_{I}]\leq 1$ for any $\mathcal{F}$-stopping time $\nu$ and for all $\mathbb{P}\in\mathcal{P}_{0}(I)$. (Ruf et al., 2023)
That is, at any $\mathcal{F}$-stopping time $\nu$, $E_{I}^{[\nu]}$ is a valid e-variable for $H_{I}$. An e-process variable is also called safe anytime-valid e-variable (c.f. Koolen and Grünwald, 2022). We denote the set of all $\mathcal{F}$-stopping times by $\mathfrak{N}$.
If $H_{I}$ is true, the probability that a corresponding e-process ever exceeds a given threshold is bounded from above. Specifically, Ruf et al. (2023) show that
$\mathbb{P}(\exists\nu\in\mathfrak{N}:E_{I}^{[\nu]}\geq\frac{1}{\alpha})\leq\alpha\quad\text{for every }\alpha\in[0,1]\text{ and every }\mathbb{P}\in\mathcal{P}_{0}(I).$ (3)
This is an adaption of Ville’s inequality. Denote by $e_{I}^{[\nu]}$ the e-process value corresponding to the e-process variable $E^{[\nu]}_{I}$. Furthermore, let $\phi_{\alpha}^{[E,\nu]}(H_{I})\coloneq\mathbb{I}\{e_{I}^{[\nu]}\geq 1/\alpha\}$ be a local level-$\alpha$ test based on the e-process value at time $\nu$ corresponding to hypothesis $H_{I}$. The probability that a true hypothesis is ever rejected by $\phi_{\alpha}^{[E,\nu]}$ is at most $\alpha$, as ensured by Eq. (3). Therefore, even though we test $H_{I}$ multiple times as more and more observations become available, $\phi_{\alpha}^{[E,\nu]}(H_{I})$ has type I error control for all $\nu\in\mathfrak{N}$. We call $\phi_{\alpha}^{[E,\nu]}$ an anytime-valid local level-$\alpha$ test at time $\nu$.
As with e-variables, there are several families of e-processes and many of them are based on a likelihood ratio (c.f. Koolen and Grünwald, 2022; Pérez-Ortiz et al., 2024; ter Schure et al., 2024; Wang and Ramdas, 2025). Alternatively, e-processes can be constructed by merging sequential (or independent) e-values (Vovk and Wang, 2024a). An overview of different families of e-processes is given in Ramdas et al. (2023).
In the next section, we combine the theory of simultaneous confidence bounds based on closed testing with the theory of e-processes and propose a procedure to compute safe anytime-valid upper confidence bounds for $\tau(R)$.
3 Anytime-valid simultaneous upper confidence bounds
We consider the filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}},\mathcal{P})$, with $\Omega$, $\mathcal{F}$ and $\mathcal{P}$ as in Section 2 and $(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}}$ the natural filtration corresponding to the observable units. That is, we assume that we sequentially observe data and at each observation time point we observe $m$ data points. We make no assumptions regarding the dependency structure of the data.
As before, we consider the set of hypotheses $\mathcal{H}$. We assume that valid e-processes for every elementary hypothesis $H_{i}$, $i\in\{1,\ldots,m\}$, are available. Furthermore, we only consider e-processes with respect to the natural filtration. Hence, the set of valid stopping times is given by $\mathfrak{N}=\mathbb{N}_{0}$ for each e-process.
Denote by $\mathcal{U}_{\alpha}^{[n]}$ the index set of the hypotheses rejected by $\phi_{\alpha}^{[E,n]}$ at time $n$. Furthermore, let $\mathcal{X}_{\alpha}^{[n]}$ be the index set corresponding to the hypotheses rejected by closed testing based on $\mathcal{U}_{\alpha}^{[n]}$. Utilizing $\phi_{\alpha}^{[E,n]}$ to determine $\mathcal{U}_{\alpha}^{[n]}$ requires valid e-processes for intersection hypotheses. It has been pointed out by Ramdas et al. (2023) that e-processes of elementary hypotheses can be merged, similarly to e-values. Lemma 1 demonstrates that $(E_{I}^{[n]})_{n\in\mathbb{N}_{0}}$ with $E_{I}^{[n]}=\sum_{i\in I}E_{i}^{[n]}/|I|$ is an e-process corresponding to the intersection hypothesis $H_{I}$.
Lemma 1.
Consider multiple e-processes with respect to $(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}}$. The average is a valid e-merging function for the e-processes and the resulting e-process is an e-process with respect to $(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}}$.
Proof.
Let $(E_{I}^{[n]})_{n\in\mathbb{N}_{0}}$ be the e-process corresponding to hypothesis $H_{I}$ with
$E_{I}^{[n]}=\sum_{i\in I}E_{i}^{[n]}/|I|\qquad\forall n\in\mathbb{N}_{0},$
where $E_{i}^{[n]}$ is the e-process variable at time point $n$ corresponding to hypothesis $H_{i}$. For any stopping time $n\in\mathbb{N}_{0}$, $\mathbb{E}_{\mathbb{P}}[E_{I}^{[n]}]=\mathbb{E}_{\mathbb{P}}[\sum_{i\in I}E_{i}^{[n]}/|I|]=\sum_{i\in I}\mathbb{E}_{\mathbb{P}}[E_{i}^{[n]}]/|I|\leq 1$ for all $\mathbb{P}\in\mathcal{P}_{0}(I)$ by definition of an e-process. Therefore, $(E_{I}^{[n]})_{n\in\mathbb{N}_{0}}$ is an e-process with respect to $(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}}$ corresponding to $H_{I}$. ∎
The following proposition is an extension of the bounds given in Eq. (1) to be anytime-valid simultaneous upper confidence bounds for $\tau(R)$.
Proposition 1.
Define
$c_{\alpha}^{[n]}(R)\coloneq\max\{|I|:I\in\mathcal{C}_{R},I\notin\mathcal{X}_{\alpha}^{[n]}\}.$
Then,
$\mathbb{P}\left(\tau(R)\leq c_{\alpha}^{[n]}(R),\qquad\forall R\in\mathcal{C},\forall n\in\mathbb{N}_{0}\right)\geq 1-\alpha,$ (4)
for all $\mathbb{P}\in\mathcal{P}_{0}(T)$ and $\alpha\in(0,1)$.
Proof.
Let $(E_{T}^{[n]})_{n\in\mathbb{N}_{0}}$ be the e-process corresponding to $H_{T}$. Because of Eq. (3), $\mathbb{P}(\exists n\in\mathbb{N}_{0}:E_{T}^{[n]}\geq 1/\alpha)\leq\alpha$ for every $\mathbb{P}\in\mathcal{P}_{0}(T)$ and every $\alpha\in[0,1]$ (Ruf et al., 2023). Thus, $\mathbb{P}(T\in\mathcal{U}_{\alpha}^{[n]})\leq\alpha$ for all $n\in\mathbb{N}_{0}$. Eq. (4) follows from the fact that $\tau(R)\leq c_{\alpha}^{[n]}(R)$ if $T_{R}\notin\mathcal{X}_{\alpha}^{[n]}$, which is the case if $T\notin\mathcal{U}_{\alpha}^{[n]}$. Since the event $T\notin\mathcal{U}_{\alpha}^{[n]}$ does not depend on the chosen discovery set $R$, the bounds are simultaneously valid for all $R\in\mathcal{C}$. ∎
Based on Proposition 1, anytime-valid simultaneous lower confidence bound for the TDP can be computed as $1-c_{\alpha}^{[n]}(R)/|R|$.
Computing $c_{\alpha}^{[n]}(R)$ at time point $n$ requires testing up to $2^{m}-1$ hypotheses, unless shortcuts are used. We propose such a shortcut when using the average as e-merging function and $\phi_{\alpha}^{[E,n]}$ to test the hypotheses. We assume that $R$ is fixed over time. The shortcut is based on the following lemma.
Lemma 2.
Denote by $K_{h}^{[n]}\in\mathcal{C}_{R}$ the set of indices of the elementary hypotheses in a rejection set $R$ with the $h$ smallest e-process values at time point $n$, $n\in\mathbb{N}_{0}$. Denote by $\bar{K}_{h}^{[n]}\notin\mathcal{C}_{R}$ the set of indices corresponding to the elementary hypotheses not in $R$ with the $h$ smallest e-process values at time point $n$. Assume that $K_{h+1}^{[n]}\in\mathcal{X}_{\alpha}^{[n]}$. Then, $K_{h}^{[n]}\notin\mathcal{X}_{\alpha}^{[n]}$ if
$\sum_{i\in K_{h}^{[n]}}e_{i}^{[n]}-\frac{h}{\alpha}<\frac{k^{\ast}}{\alpha}-\sum_{j\in\bar{K}_{k^{\ast}}^{[n]}}e_{j}^{[n]},$ (5)
where $k^{\ast}=argmax\{k/\alpha-\sum_{j\in\bar{K}_{k}^{[n]}}e^{[n]}_{j}\}$, $0\leq k\leq m-|R|$, with $\sum_{j\in\bar{K}_{0}^{[n]}}e^{[n]}_{j}\coloneq 0$ by convention.
Proof.
Remember that $K_{h}^{[n]}\notin\mathcal{X}_{\alpha}^{[n]}$ if there exists at least one $J\supseteq K_{h}^{[n]}$ such that $J\notin\mathcal{U}_{\alpha}^{[n]}$. Let $J\coloneq\{K_{h}^{[n]},\bar{K}_{k^{\ast}}^{[n]}\}$, where $\bar{K}_{k^{\ast}}^{[n]}=\emptyset$ is allowed. Since $K_{h+1}^{[n]}\in\mathcal{X}_{\alpha}^{[n]}$, we do not consider any $J$ which includes additional rejected hypotheses, as we know that this hypothesis is rejected by closed testing. When using the local level-$\alpha$ e-test $\phi_{\alpha}^{[E,n]}$ and the average as an e-merging function, $J\notin\mathcal{U}_{\alpha}^{[n]}$ if
$\frac{\sum_{i\in K_{h}^{[n]}}e_{i}^{[n]}+\sum_{j\in\bar{K}_{k^{\ast}}^{[n]}}e_{j}^{[n]}}{h+k^{\ast}}$ $<\frac{1}{\alpha}$
$\Leftrightarrow\sum_{i\in K_{h}^{[n]}}e_{i}^{[n]}-\frac{h}{\alpha}<\frac{k^{\ast}}{\alpha}-\sum_{j\in\bar{K}_{k^{\ast}}^{[n]}}e_{j}^{[n]}.$
Let $\mathcal{I}_{h}\subseteq\mathcal{C}_{R}$ be the index set of the rejected intersection hypotheses of size $h$, $h\in\{1,\ldots,|R|\}$. Lemma 2 allows us to sequentially test if there exists an hypothesis $H_{I_{h}}$ with $I_{h}\in\mathcal{I}_{h}$ for which Eq. (5) holds at time point $n$, with $h$ iteratively decreasing until $I_{h}\notin\mathcal{X}_{\alpha}^{[n]}$ and then $c_{\alpha}^{[n]}(R)=h$. Therefore, for a fixed rejection set $R$, we need to test at most $|R|$ hypotheses at time point $n$.
Furthermore, notice that the right side of Eq. (5) does not depend on $h$ and can be viewed as a rejection set specific threshold for testing any hypothesis $H_{I}$, $I\in\mathcal{C}_{R}$ at time $n$.
4 Practical considerations
In this section, we consider the implementation of the proposed anytime-valid simultaneous lower confidence bound for the TDP. We first comment on the behavior of the confidence bounds over time. As more observations become available, the bounds should become more informative, i.e., they should become closer to the true TDP. We say that the bounds are tight if they are close to the true TDP. The lower confidence bounds bounds become tighter over time if they are non-decreasing. Another important aspect in the practical application of the proposed procedure is the choice of the e-process used to define the local level-$\alpha$ test. We comment on that in the second part of this section.
4.1 Non-decreasing anytime-valid simultaneous confidence bounds
Whether the proposed lower confidence bounds for the TDP are non-decreasing or not depends solely on the utilized e-process. While the probability of the e-process crossing a given threshold is bounded from above if the corresponding hypothesis is true, there are no guarantees that the e-process does not cross a given threshold multiple times if the hypothesis is false. That is, if $H_{I}$ is false, $\phi_{\alpha}^{[E,n]}$ might reject $H_{I}$ at a given time point $n$, but $\phi_{\alpha}^{[E,(n+k)]}$ might not reject (”accept”) $H_{I}$ at time point $n+k$, $k\in\mathbb{N}_{0}$. However, for the upper confidence bound to be non-decreasing, we require that only accept-to-reject decisions (ARD) are made, i.e., once an hypothesis is rejected, it remains rejected.
In this work, we consider two approaches to ensure ARD. First, the bounds defined by Eq. (1) with the local level-$\alpha$ test based on an e-to-p process for the elementary hypotheses can be used. An e-to-p process for hypothesis $H_{i}$ is a p-process which is defined as $p_{i}^{[n]}\coloneq\min_{\ell\leq n}\{1/e_{i}^{[\ell]}\}$ (Ruf et al., 2023). To compute a p-process for an intersection hypothesis $H_{I}$, the e-to-p processes corresponding to all $H_{i}$, $i\in I$, are merged at each time point as described in Goeman et al. (2019). Subsequently, the computational shortcut to reduce the number of effectively tested hypotheses as introduced by Goeman et al. (2019) can be applied. However, when ARD is ensured through the e-to-p process, the confidence bounds are not valid under arbitrary dependence of the $m$ p-values at time $n$. For more information, see Goeman et al. (2019) and references therein.
Second, ARD can be ensured by defining the local level-$\alpha$ test based on the maximum value of the so-far observed e-process, that is, $\phi_{\alpha}^{[E,n_{\ell}]}(H_{I})\coloneq\mathbb{I}(\max_{\ell\leq n}e_{I}^{[\ell]}\geq 1/\alpha)$. While $\max_{\ell\leq n}E_{I}^{[\ell]}$ itself is not a valid e-process, $\phi_{\alpha}^{[E,n_{\ell}]}$ is a valid local level-$\alpha$ test due to Eq. (3). Alternatively, one could also define the local level-$\alpha$ test as $\phi_{\alpha}^{[E,\ddot{n}]}\coloneq\mathbb{I}(\ddot{e}_{I}^{[n]}\geq 1/\alpha)$, with
$\ddot{e}_{I}^{[n]}\coloneq\begin{cases}e_{I}^{[n]}\quad n<\ddot{n}\\ e_{I}^{[\ddot{n}]}\quad n\geq\ddot{n},\end{cases}$
where $\ddot{n}\coloneq\min\{n|\phi_{\alpha}^{[E,n]}(H_{I})=1\}$ is the time point at which $H_{I}$ is rejected by the local test for the first time. Using $\phi_{\alpha}^{[E,\ddot{n}]}$ reduces the computational effort compared to $\phi_{\alpha}^{[E,n_{\ell}]}$ or $\phi_{\alpha}^{[E,n]}$, because once $\ddot{e}_{I}^{[n]}\geq 1/\alpha$, it does not need to be updated. We want to emphasize that these ARD insurances still require the computation of valid e-processes for all elementary hypotheses at every time point, even after a local level-$\alpha$ test has rejected some elementary hypotheses. This is because computing the e-process corresponding to a (not yet rejected) intersection hypothesis using the average requires valid e-processes for the corresponding elementary hypotheses.
4.2 Choice of the e-process
The behavior of the proposed procedure depends on the choice of the e-process defining the local level-$\alpha$ test. There are four main aspects which need to be considered when choosing the e-process.
First, the faster the utilized e-process can gather evidence against the null hypothesis, the faster the confidence bounds become tight. Therefore, we aim to use an e-process which gathers evidence as fast as possible. The GROW criterion can be applied to e-processes as well. However, while e-processes fulfilling the GROW criterion have the optimal growth-rate in the worst case scenario, they might have a slower growth-rate than other e-processes if the worst-case scenario does not occur. Therefore, in practice an e-process which has a comparatively large growth-rate across multiple scenarios is desirable.
Second, the proposed procedure is valid as long as the utilized e-process is valid. Thus, the dependency structure as well as the distribution of the data need to be considered. This includes considering the available information about the distribution of the data, for example, if the variance-covariance matrix is known or not. This is similar to choosing an appropriate test statistic when working with p-values.
Third, the set of possible stopping times of the utilized e-process needs to be taken into account. We require that the the e-process is defined on the natural filtration. However, another filtration (coarser or finer than the natural filtration) might be more practical depending on the application.
Lastly, the computational effort regarding e-processes is of interest in practice. Computing anytime-valid simultaneous lower confidence bounds for the TDP requires computing e-processes for all of the $m$ elementary hypotheses. The number of considered hypotheses $m$ can be very large. For example, in fMRI more than 100.000 hypotheses are of interest. Hence, the computational effort, and therefore the time needed to compute a single e-process, should be reasonably small.
An in-depth comparison of e-processes regarding these four aspects is outside the scope of this paper. Ly et al. (2024) compare four e-processes in terms of their growth-rate in different scenarios. These e-processes correspond to the likelihood ratio of an t-test statistic, a test statistic often applied in practice. All four compared e-processes are based on the work by Gronau et al. (2020) and Pérez-Ortiz et al. (2024). The family of e-processes is defined by
$lr_{n}(t)\coloneq\frac{\mathcal{L}(t|\sqrt{n_{t}}\delta,\lambda)}{\mathcal{L}(t|0,\lambda)},$ (6)
where $\mathcal{L}(\cdot|\mu,\lambda)$ is the likelihood of the t-distribution with $\lambda$ degrees of freedom and non-centrality parameter $\mu$. Here, $n_{t}$ is the effective sample size used to compute the t-test statistic $t$. Furthermore, $\delta$ is a parameter that is chosen such that $lr_{n}(t)$ is an e-process. There are different ways to define $\delta$. Ly et al. (2024) recommend using a non-local moment prior on $\delta$, which is a two-bump prior with bumps at $\delta_{min}$ and $\delta_{min}$, where $\delta_{min}$ is the minimal relevant effect size. They name the resulting e-process the ”mom” e-process. While the mom e-process does not necessarily have the largest growth rate in all scenarios, it’s growth-rate appears to be stable across different scenarios.
5 Simulations
In this section, we investigate the behavior of the proposed procedure in a simulation study. In our simulated experiment, $m=90$ hypotheses have been tested based on the data of up to $N=100$ subjects. We simulate that we have sequentially observed data, with data from one subject at each time point. This is in accordance with fMRI studies, where subjects need to be scanned one after the other. We have tested whether the mean is larger than zero, i.e., $H_{1}:\mu_{1}\leq 0,\ldots,H_{m}:\mu_{n}\leq 0$. Half of the hypotheses have been true. For each subject, the data are identically distributed, i.e., $\boldsymbol{y}_{n}=(y_{\ell,1},\ldots,y_{\ell,m})^{\top}$, with $\boldsymbol{y}_{n}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{V})$ for each $n=1,\ldots,N$ and
$\boldsymbol{\mu}$ $=(0,\ldots,0,\mu_{alt}\ldots,\mu_{alt})^{\top},$
$\boldsymbol{V}[a,b]$ $=\begin{cases}1\qquad\text{if }a=b,\\ \rho\qquad\text{else.}\end{cases}$
We have set $\mu_{alt}\in\{0.5,1,1.5\}$ and $\rho\in\{0.2,0.6\}$.
The e-processes corresponding to the elementary hypotheses have been mom e-processes. The e-processes have been computed utilizing the R package ”safestats” (Turner and Ly, 2022), with $\delta_{min}=0.5$. The closed testing procedure has been based on $\phi_{\alpha}^{[E,n]}$ as well as $\phi_{\alpha}^{[E,\ddot{n}]}$. We say that the anytime-valid simultaneous lower confidence bounds for the TDP are based on the mom e-process and the mom_ARD e-process when we have utilized $\phi_{\alpha}^{[E,n]}$ and $\phi_{\alpha}^{[E,\ddot{n}]}$, respectively. To ensure ARD we have also computed the e-to-p process based on the mom e-process. To compute the confidence bounds at each observation time point, we have used the procedure as implemented in the package R ”hommel” (Goeman et al., 2021a) and described by Goeman et al. (2019). We say that the corresponding confidence bounds are based on the e-to-p process.
We have considered nine different rejection sets with TDP of $\pi_{1}\in\{0.1,0.2,\ldots,0.9\}$. For each rejection set we have computed the $1-\alpha$ anytime-valid simultaneous lower confidence bounds for the TDP, $\hat{\pi}_{1}$, based on the three processes, with $\alpha=0.2$. The bounds are based on at least 11 observations. The results are based on 1000 Monte-Carlo iterations.
We examine the validity of the bounds first. If the proportion of iterations in which $\hat{\pi}_{1}>\pi_{1}$ is at most $\alpha$, the bounds are considered valid. At every observation time point the proposed procedure has returned valid bounds, regardless of the process it is based on, the covariance $\rho$ or the effect size $\mu_{alt}$. Figure 1 displays the proportion of iterations in which $\hat{\pi}_{1}>\pi_{1}$ over time for $\rho=0.2$, $\pi_{1}\in\{0.1,0.9\}$ and $\mu_{alt}\in\{0.5,1,1.5\}$. For small $\pi_{1}$, the proportion of iterations with $\hat{\pi}_{1}>\pi_{1}$ has been larger than for large $\pi_{1}$. Furthermore, the proportion of iterations with $\hat{\pi}_{1}>\pi_{1}$ has been lowest for the proposed procedure based on the mom process. This is to be expected, since utilizing the mom process can lead to decreasing bounds. When increasing the effect size $\mu_{alt}$, the maximum of the proportion of iterations with $\hat{\pi}_{1}>\pi_{1}$ has increased as well. The results for $\rho=0.6$ have been similar. Under ARD insurance, the proportion of iterations in which $\hat{\pi}_{1}>\pi_{1}$ is generally slightly larger for $\rho=0.6$ compared to $\rho=0.2$.
Figure 1: Empirical non-coverage rate over time for $\rho=0.2$. The three different utilized local level-$\alpha$ e-tests are based on the mom e-process values (mom), the mom e-process values adapted to only allow accept-to-reject decisions (mom_ARD) and the e-to-p process values based on the mom e-process (e-to-p), respectively. The black solid line corresponds to the significance level $\alpha=0.2$. Results are based on 10001000 iterations.
Next, we investigate the power of the proposed procedure. The procedure is considered to be powerful when the lower confidence bound $\hat{\pi}_{1}$ is large. Figure 2 displays the changes in the average of the confidence bounds across the iterations as the number of available observations increases for $\pi_{1}\in\{0.1,0.9\}$, $\mu_{alt}\in\{0.5,1,1.5\}$ and $\rho=0.2$. Generally, the proposed procedure based on the e-to-p process has had a more gentle growth rate over the first few observations compared to the proposed procedure based on the mom and mom_ARD e-process. On average, the bounds based on all three processes have taken a similar time to converge to the true TDP. When increasing the effect size $\mu_{alt}$, the time to converge has decreased. Utilizing the mom e-process has returned looser bounds compared to using the e-process adapted to ARD. The difference between the average TDP bounds based on different e-processes has been larger for small $\pi_{1}$ than for large $\pi_{1}$. Once again, results for $\rho=0.6$ have been similar. As before, the bounds under ARD insurance have been slightly more powerful for $\rho=0.6$ compared to $\rho=0.2$.
Overall, these simulations indicate that the proposed procedure returns valid lower confidence bounds for the TDP. For a reasonably large effect size ($\mu_{alt}\geq 1$), the bounds seem to converge to the true value of the TDP in a reasonable time, in our simulations after roughly 30 observations or less.
Figure 2: Average of the anytime-valid simultaneous lower confidence bound $\hat{\pi}_{1}$ for the true discovery proportion $\pi_{1}$ over time for $\rho=0.2$. The three different utilized local level-$\alpha$ e-tests are based on the mom e-process values (mom), the mom e-process values adapted to only allow accept-to-reject decisions (mom_ARD) and the e-to-p process values based on the mom e-process (e-to-p), respectively. The black solid line corresponds to $\pi_{1}$. Results are based on 10001000 iterations.
6 Discussion
We have proposed a novel procedure to compute anytime-valid simultaneous lower confidence bounds for the TDP. To this end, we have utilized confidence bounds based on closed testing, where the closed testing rejection sets are determine by local tests defined by e-processes. To make the computation of the proposed bounds feasible in settings where the number of considered hypotheses is large, we have derived a computational short-cut.
In contrast to established procedures, the proposed method allows for optional stopping. That is, the data gathering process may be stopped after any observational time point and the decision to continue may rely on the data observed so far. Optional stopping is a desirable property in applications where data acquisition is both time and cost intensive. The proposed method can be applied in research areas such as genomics and neuroscience.
The presented procedure has been derived under the assumption that the utilized e-processes are defined with respect to the natural filtration. While this seems to be a sensible requirement in practical settings, it is not a necessary requirement. Indeed, the procedure can be applied using e-processes for any other filtration $(\mathcal{G}_{n})_{n\geq 0}$ as well. In this case, the procedure only returns valid results for $\mathcal{G}$-stopping times.
We have commented on practical considerations, namely on ensuring that the bounds are non-decreasing and on the choice of the underlying e-process. By choosing a suitable e-process, the method can be applied to any temporal dependency structure. Regardless of the utilized e-processes, the proposed procedure allows for arbitrary dependence between the considered hypotheses.
We have studied the proposed method’s behavior using simulations. In the simulations, the confidence bounds converged to the true TDP as the number of observations increased. The time to converge depended on the distribution of the data, the bounds converged faster to the true TDP if the effect size and/or the covariance increased. Furthermore, results of the simulation study suggest that utilizing e-processes with ARD insurance returns the most powerful bounds.
Subsequent work entails specifying guidelines for the application of the proposed procedure, focusing on the choice of the e-process used in the method. We aim to apply the proposed procedure to real data in the future as well, such as data of an fMRI experiment.
7 Acknowledgement
Friederike Preusse gratefully acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)- project number 281474342.